A fair die is rolled 8 times. What is the probability that an even number (2,4, 6) will occur between 2 and 4 times? Round your answer to four places after the decimal
Answer by stanbon(75887)

A fair die is rolled 8 times. What is the probability that an even number (2,4, 6) will occur between 2 and 4 times? Round your answer to four places after the decimal
--------
Binomial Problem with n = 8 ; p(even) = 1/2 ; P(odd) = 1/2
--------
P(2<= x <=4) = binomcdf(8,1/2,4)-binomcdf(8,1/2,1) = 0.6016
====================
Cheers,
Stan H.